Explain the meaning of a degenerate gas. A white dwarf star has a mass of 1030kg. Its luminosity is 1024 Js^−1. Calculate how long it can survive with its present luminosity if its internal temperature is 107K.
1
Expert's answer
2020-03-10T11:31:04-0400

Using equation we can find time


"t=10^5\cdot ((\frac{L_s}{L})^3(\frac{M}{M_s})^4)^{\frac{1}{5}}"

"t=10^5\cdot ((\frac{3.828\cdot10^{26}}{10^{24}})^3(\frac{10^{30}}{2\cdot10^{30}})^4)^{\frac{1}{5}}=2.1\cdot10^6 years"